### v3

# Supplemental Material Appendix A Additional Related Works A.1 Image Denoising and Restoration

_main
[…]
Supplemental Material
Appendix A Additional Related Works
A.1 Image Denoising and Restoration
Image denoising and restoration belong to a set of problems called image inverse problems [Ongie2020-zi]. There are several different levels of difficulty for image denoising and restoration problems. The easiest level is when the corruption process is known as a linear process (such as blurring, inpainting, uniform Gaussian noise, etc) and can be programmatically replicated onto clean images, and the goal is to create an inverse process to the corruption. There have been extensive studies of restoring known corruptions, including supervised approaches [Ongie2020-zi] and unsupervised approaches [Kawar2022-vc, Kadkhodaie2021-wd, wang2022zero]. The second level of difficulty is when the corruption process is not reproducible, but there are existing pairs of clean and corrupted images allowing for supervised training, and the typical solution is training models to generate the clean image given the corrupted images in a supervised way, such as UNet [Ronneberger2015-um, Manifold2019-ft] or Conditional Diffusion [Choi2021-mx, Kawar2022-vc, Saharia2022-co]. The most difficult setting is when we don’t have paired data and also cannot reproduce the corruption programmatically; instead, we only have access to unpaired clean images and corrupted images. Existing approaches for this challenging setting include cycle-consistency-based approaches such as CycleGAN [Zhu2017-wf] and UnitDDPM [Sasaki2021-ks], VAE-based approaches [Zheng2023-dz], and flow-based approaches [Du2020-om]. In addition to the above settings, there are also denoising methods that do not require a training dataset, such as MedianBlur or Deep Image Prior [Ulyanov2016-kz]. General Diffusion Prior [fei2023generative] is a recent unsupervised diffusion-based method that can perform unpaired non-linear blind image restoration, but it is extremely computationally costly as it requires full generative diffusion with multiple gradient steps within each reverse diffusion step.
Restoring noisy SRH images generally falls within the third difficulty level (i.e. no paired data and not a linear inverse problem), as the distribution of noise is unpredictable. Although it is possible to apply supervised methods from the second level by training to remove Gaussian and Laser noises (such as our UNet and Conditional Diffusion baselines), they are generally outperformed by unpaired methods as shown in our experiments. Several recent approaches (including [chung2022come], [Chung2023-qt], and [he2023tdiffde]) also used the last part of the reverse diffusion process for unpaired image restoration. We discussed these closely related works in more detail in Section LABEL:sec:relwork.
A.2 Deep Learning applications on SRH
Deep learning techniques have been playing an important role ever since SRH was used for intra-operative neurosurgery, as the very first paper that introduces this optical diagnostic technique [Orringer2017-nn] trained a multi-layer perceptron to automatically predicts brain tumor subtypes from quantified SRH image attributes. Later, Hollon et al. [Hollon2020-ez] proposed a CNN-based model that can directly interpret raw SRH images (instead of quantified attributes) and make diagnosis at over 94% accuracy. The first public dataset of clinical SRH images, OpenSRH [Jiang2022-cj], was released in 2022, and it contains SRH images from 300+ brain tumor patients and 1300+ unique whole slide optical images. DeepGlioma [Hollon2023-uf] has allowed deep-learning-based models to perform molecular classification of diffuse glioma types via a multimodal model that takes in both SRH images and diffuse glioma genomic data. HiDisc [Jiang2023-rq] and S3L [hou2024selfsupervised] are self-supervised visual representation learning methods that were developed to classify whole-slide SRH images.
All the above works focused on the classification of SRH images, and there have been few previous works that attempted to tackle the unavoidable noise problem of SRH images through image restoration. To the best of our knowledge, the only previous work on this topic was [Manifold2019-ft], in which the authors collected a paired dataset using low and high laser energy, and trained a U-Net to perform the denoising. In comparison, our method works for all types of noisy SRH images (not just those resulting from low laser power), and we have shown that our method outperforms their U-Net approach.
Appendix B Details about SRH images
We include examples of image degradation for different cytologic/histologic features and brain tumor types in Figure 1. The examples show there exist many vastly different types of degradation that can randomly occur in low-quality SRH images, therefore making SRH image restoration a challenging task, but RSCD was able to restore them quite well.
[width=]figs/Fig_10.pdf
Appendix C Methodology Additional Details
C.1 Step Calibrator Training and Calibration Distribution
We show the full algorithm of training the step calibrator with manually added Gaussian noise and augmentations in Algorithm C.1, as well as one example of the augmentation in Figure 2.
[…]
To determine a proper , we ran our trained step calibrator on 63K real low-quality SRH images, and the distribution of is shown in Figure 3. We can see that the number of images decreases exponentially with its , and never exceeded 200, which justifies our choice of in our diffusion model training.
[width=0.6]figs/aug.png
[width=0.6]figs/histogram.png
C.2 Diffusion Model Training
We show the full training algorithm in Algorithm C.2 below. It is almost identical to the training algorithm from [Ho2020-tm], except that we only sample up to .
[…]
{algorithmic}[1]
[…]
Appendix D Experiment details and additional results
D.1 Baseline and Ablation details
Below are the modeling and training details of the baselines:
CycleGAN [Zhu2017-wf]: We trained a CycleGAN model for image-to-image translation between unpaired low-quality and high-quality images. We used the 4.5K hand-picked high-quality and 3.8K hand-picked low-quality images for training the CyleGAN. Both the high-quality and low-quality image sets contained at least 10 different tumor types. The CycleGAN model performed better when trained on the manually selected images compared to training on the full 840K filtered high-quality images.
Synthetic noise[Lehtinen2018-zd, Manifold2019-ft]: A U-Net model is trained by adding Gaussian noise to high-quality images to generate synthetic low-quality images. We train only with the 4.5K manually selected high-quality images because we found that the model performs better compared to training on the full 840K filtered high-quality images.
Conditional Diffusion [Kawar2022-vc, Saharia2022-co]: We train a conditional generative diffusion model, where we sample from a random Gaussian prior, and generate the restored image conditioned on the low-quality image. The conditioning method is concatenation (i.e. the input image and the conditioning image is concatenated as input to the UNet). We train the model with the 840K HQ training images and fine-tune it with the 4.5K hand-picked images. During training, we add a random amount of Gaussian noise to the high-quality images as conditioning images. We show the full training algorithm in Algorithm.
Conditional Diffusion: Training
{algorithmic}[1]
[…]
Regularized Reverse Diffusion (RRD) [Chung2023-qt] is a recent method developed for restoring noisy medical images. It also uses part of the reverse diffusion process as restoration, but it employs a non-parametric noise estimator [Chen2015-ij] to determine the step numbers, and uses a Fourier-feature-based regularization to reduce hallucination. Since we were unable to find any official code for the RRD paper and we deal with images in a vastly different domain from theirs (RRD was originally designed for denoising single-channel MRI images), we trained our own diffusion model following the same process as [Ho2020-tm] and using the same cosine noise schedule as RSCD.
Deep Image Prior [Ulyanov2016-kz] is an unsupervised image restoration method that works by training a ResNet to reconstruct the noisy image but takes the generation of the model before its convergence as the denoised image. We take the output of step 1800 as our baseline, as images generated from this training step tend to have the best quality.
Median Blur is a traditional, but effective, method to smooth noisy images. It replaces each pixel value with the median of the pixel kernel ( for experiments).
Below are details of the ablations:
No Dynamic Recalibration: RSCD without Dynamic Recalibration, i.e. Algorithm LABEL:alg:dr without the blue parts. We take the initial step calibration and directly run that many steps of denoising to the end.
Nonparametric Noise Estimator: We replace our ResNet-based step calibrator with the nonparametric noise estimator [Chen2015-ij] used in RRD [Chung2023-qt]. We determine the step number as the closest cumulative noise to the predicted noise level from the nonparametric noise estimator.
No Step Calibrator, 10 steps always: Instead of using a step calibrator and Dynamic Recalibration, we always only perform 10 steps of diffusion-based denoising on the low-quality images.
No Step Calibrator, 50 steps always: Instead of using a step calibrator and Dynamic Recalibration, we always only perform 50 steps of diffusion-based denoising on the low-quality images.
Linear noise schedule: We use a linear noise schedule instead of the cosine linear schedule.
No Augmentation Step Calibrator: We use a step calibrator that is trained without augmentations (i.e. Algorithm C.1 without the red parts).
D.2 Unpaired Images Evaluation additional qualitative results
We show additional qualitative comparisons of restorations of unpaired images in Figure 4.
[width=]figs/supp.unpaired.pdf
D.3 Near-registered dataset collection details
Imaging with low laser power is one of the known ways to produce noisy images. We generate a paired set of near-registered SRH images by imaging a surgical specimen immediately after the SRH imaging system is started (without laser warm-up and optimization), and then image the same specimen after 20 minutes, once the laser has been optimized and sufficiently warmed up. Next, image registration is required to spatially align the low-quality and high-quality images because the tissue specimen may be shifted due to gravity between the two image acquisitions. Paired images were registered by fitting an affine transformation using scale-invariant feature transform [Lowe2004-gq] and RANSAC [derpanis2010overview].
While the near-registered images are matched enough for meaningful evaluation of fidelity, they are not fitting for training supervised image restoration because (1) the images are not perfectly paired (free-floating structures like red blood cells may have rotated/shifted relative to nearby cells), so training with them is just promoting hallucinations; (2) the noise in this dataset results from cold laser alone, so it does not well represent all types of random noise that could occur in SRH images from real intraoperative settings.
D.4 Hyperparameters
We list all architecture/training/inference hyperparameters for all components of RSCD in Table 1.
[…]
Base Model | ResNet-50 | |||
Prediction Head |
Learning Rate | 0.00001 | ||||
Optimizer | Adam | ||||
Epochs | 500 | ||||
Batch Size | 150 | ||||
[…]
Model | UNet | |||
In channels | 2 | ||||
Init dim | 128 | ||||
Out channels | 2 | ||||
Dim mult | [1,2,2,4] | ||||
Resnet block groups | 8 | ||||
Convnext Mult | 2 | ||||
[…]
Noise Schedule | Cosine (s=0.008) | |||
Total Steps (T) | 1000 | ||||
200 | ||||
Batch Size | 8 | ||||
Learning Rate | 0.00001 | ||||
Loss type | L1 | ||||
Epochs |
Optimizer | Adam | ||||
[…]
D.5 Computing Resource and Inference Efficiency
Training the step calibrator takes about 5 hours on 1 A40 GPU. Training the diffusion model on 840K data took about 4 days on 1 A40 GPU, and fine-tuning on 4.5K hand-picked data took about 1 day on 1 A40 GPU.
We analyze the inference efficiency of RSCD on a single RTX 2080 GPU, in both runtime and NFE of the forward function of the UNet. On average over 1000 randomly selected low-quality SRH images, RSCD takes about 14.5 seconds to denoise a 256x256 patch (34 NFE), while methods that runs full generative reverse diffusion (e.g. Conditional Diffusion) take 431 seconds per patch (1000 NFE). Thus, RSCD is about 30x faster than running a full generative reverse diffusion process. Also, note that this number is obtained by denoising each patch individually without batching on an old and slow GPU for fair comparison. The amortized computation time can be improved by a lot if we perform the denoising steps in large batches and with a better GPU.
D.6 Human Expert Preference Details
We show the human expert preference collection interface in Figure 5. The expert is given the original LQ image as well as 2 restorations (one from RSCD and one from a baseline, in random order; the expert is not told which restoration is baseline and which one is RSCD), and then the expert is asked to (1) select one restoration that the expert believes has better overall quality, and (2) determine which restoration has clinically significant hallucinations (can choose ”neither” if neither restoration hallucinates). We perform a total of 1200 such pair-wise preference collections (100 LQ images x 6 baseline comparisons x 2 preferences from different experts per comparison), distributed among all participating experts.
[width=]figs/interface.pdf
Appendix E Downstream task details
E.1 Deep learning-based tumor diagnosis
Data:
A portion of our SRH images have ground truth labels on whether the tissue is normal or tumor. Therefore, we took the intersection of that labeled portion and the 12K unpaired low-quality images used for the experiment in section LABEL:sec:unpaired as the dataset for deep learning-based tumor diagnosis. The dataset contains 8,372 low-quality images, each with its ground truth label of tumor/normal.
Model:
We used the segmentation model from [Hollon2020-ez], the most influential and widely accepted study on deep learning-based brain tumor diagnosis through SRH images, as our classification model. The model classifies each SRH image into one of three classes: normal, tumor, and non-diagnostic. The model was trained to classify an SRH image as non-diagnostic if the model is unable to extract sufficient information from the image to make a diagnosis, and there are several possibilities for an image to be non-diagnostic, including: (1) the image scanning an irrelevant part of the slide, or (2) the image quality is too low.
Evaluation:
For each image in our dataset, we run the same classification model on both the original and the restored version of the image. We categorize the predictions into 3 categories as follows:
Correct: If the ground truth label matches the predicted label (i.e. predicting an image from a tumor tissue as tumor, or predicting an image from a normal tissue as normal)
Wrong: If the ground truth label is the opposite of the predicted label (i.e. predicting an image from a tumor tissue as normal, or predicting an image from a normal tissue as tumor)
Non-Diagnostic: If the predicted label is non-diagnostic
As shown in Table LABEL:fig:classifier, the restored images have a significantly higher “correct” rate while having a much lower percentage of misclassified or non-diagnostic images. This means that RSCD can restore a significant portion of the images that were originally classified as non-diagnostic or wrongly classified due to poor quality, and make them high-quality enough to be correctly classified.
We include examples of original/restored images with their automated diagnosis results in Figure 6.
[width=]figs/Fig_15.pdf
E.2 Z-stack Image Restoration
Image acquisition.
Non-destructive z-stack SRH imaging often results in noisy images in deeper slices due to tissue scattering from the structures in the slices above, especially in specimens that contain blood or necrotic tissue. Z-stack SRH imaging follows the protocol for normal SRH imaging. Once the scanning of a whole slide is complete, the SRH imaging system automatically adjusts the laser focus, and an image at a higher depth is subsequently acquired. The z-stacked images are acquired starting at the default depth of 22 , with a z-resolution of 2 for 20 z-sections (reaching a final depth of 60 ). Each z-stacked whole-slide is then patched into volumes for denoising.
Additional results.
We include additional examples of restored z-stack images spanning a variety of tumor types in Figure 7.
[width=.9]figs/supp.downstream.zstack.pdf